OpenStudy (anonymous):

Evaluate the following expression. 6^-3. Needs to be in a fraction please help.

jimthompson5910 (jim_thompson5910):

use the rule that \[\large x^{-k} = \frac{1}{x^k}\] to get \[\large 6^{-3} = \frac{1}{6^3} = ???\]

OpenStudy (anonymous):

in other words, whenever a number is raised to a negative number, you put a 1 over it and make the number it's being raised to positive. so, 2^-3 is 1/2^3. 6^-20 is 1/6^20, etc.